how many electric lines passing in 2c electric field (beta = 1/2)?

Grade:12

1 Answers

Mohd Shahnawaz
5 years ago
here electric flux =q/Eo  so q =2c Eo = 8.85×10-12 so the ans is 2.something into 1011electric field of lines